# OM2P: Offline Multi-Agent Mean-Flow Policy

### v1

@@ 5.2 Empirical Results
5.3 Ablation Study
We perform ablation studies to investigate two key aspects: the sensitivity of critical hyperparameters and the impact of core components within OM2P. First, we examine the role of the mixing coefficient $\eta$, which balances the behavior cloning (BC) loss and the Q-learning objective. Figure 3 presents the performance of our method across four datasets in the MPE World task under varying $\eta$ values. The results indicate that $\eta$ is a crucial hyperparameter that must be tuned in accordance with the quality of the offline dataset to achieve optimal results. Notably, small $\eta$ values lead to suboptimal learning due to insufficient Q-value guidance, while moderate to large values (e.g., $10^{3}$) consistently deliver stronger performance, particularly on high-quality datasets. This underscores the importance of properly calibrating $\eta$ to strike a balance between policy imitation and reward-driven optimization.

### v2

@@ 5.3 Ablation Study
Hyperparameters.
We examine the role of the mixing coefficient $\eta$, which balances the behavior cloning (BC) loss and the Q-learning objective. Figure 3 presents the performance of our method across four datasets in the MPE World task under varying $\eta$ values. The results indicate that $\eta$ is a crucial hyperparameter that must be tuned in accordance with the quality of the offline dataset to achieve optimal results. Notably, small $\eta$ values lead to suboptimal learning due to insufficient Q-value guidance, while moderate to large values (e.g., $10^{3}$) consistently deliver stronger performance, particularly on high-quality datasets. This underscores the importance of properly calibrating $\eta$ to strike a balance between policy imitation and reward-driven optimization.
Ablation on different components.
@@
Robustness across dataset scales.
To further evaluate the robustness of our method, we conduct ablation studies on different sizes of offline datasets under the World task expert dataset. As shown in Table 6, OM2P consistently maintains high performance even when the dataset size is significantly reduced, demonstrating strong robustness and data efficiency in offline MARL.
Table 6: Performance of OM2P under different dataset sizes on the World task expert dataset.
Number of Trajectories 4e4 3.5e4 3e4
[…]
The ablation results highlight the effectiveness and robustness of OM2P. Proper tuning of the mixing coefficient $\eta$ is essential to balance policy imitation and reward guidance, with moderate to large values yielding consistently strong performance. Modular ablations confirm that each core component—Q-function supervision, behavior cloning, and generalized timestep weighting—contributes substantially to overall performance, and removing any element degrades returns. Furthermore, OM2P maintains high performance even under reduced dataset sizes, demonstrating strong data efficiency and robustness. Collectively, these findings validate the design choices of OM2P and underscore its ability to deliver stable, efficient, and scalable offline multi-agent reinforcement learning.
@@ 6 Conclusion
7 Acknowledgement
This work is supported by the National Natural Science Foundation of China Grant 52494974.